# Trans-dimensional Hamiltonian model selection and parameter estimation from sparse, noisy data

### v1

@@ 3.4 Hybrid algorithm
Hybridizing Markov Chain Monte Carlo (MCMC) algorithms can enhance sampling efficiency and model selection for complex inverse problems. A robust approach combines the Random Walk Metropolis-Hastings (RWMH) algorithm for exploring continuous parameter spaces with Reversible Jump MCMC (RJMCMC) for transdimensional moves that enable discrete model selection. To improve within-model fits, particularly in highly non-linear and non-convex likelihood landscapes, parallel tempering (PT) based on RWMH can be incorporated for a fixed number of discrete parameters, allowing for enhanced exploration of local minima and better posterior sampling. This multi-tiered strategy balances global model selection with efficient local exploration, mitigating issues like poor mixing and mode trapping, which are common in high-dimensional and multi-modal inverse problems.
@@ 4.1 Hamiltonian and forward models
$\mathcal{H}_{k}=D\mathcal{S}_{z}^{2}+\gamma_{e}B_{z}\mathcal{S}_{z}+\sum_{i=1}^{k}\gamma_{n}B_{z}\mathcal{S}_{I_{z,i}}+\sum_{i=1}^{k}\mathcal{S}_{z}(A_{\parallel,i}\mathcal{I}_{z,i}+A_{\perp,i}\mathcal{I}_{x,i})$ (3)
where the known quantities are $D=2\pi\times 2.87$ GHz, $B_{z}$ is the strength of an externally applied magnetic field along the axis of the NV center, $\gamma_{e}$ is the gyromagnetic ratio of an electron, $\gamma_{n}$ is the nuclear gyromagnetic ratio of the nuclei, and $\mathbf{\mathcal{S}}=(\mathcal{S}_{x},\mathcal{S}_{y},\mathcal{S}_{z})$ and $\mathbf{\mathcal{I}}=(\mathcal{I}_{x},\mathcal{I}_{y},\mathcal{I}_{z})$ are the spin operators for the electronic and nuclear spins, respectively. We neglect nuclear–nuclear interactions in the Hamiltonian, as we expect these interactions to be negligible for the relatively short $\tau$ (< 10 $\mu$s) and relatively low pulse number ($\leq$ 32 CP pulses) in the experiments we consider for high-throughput characterization. We seek to recover the number of spins $k$, which determines the dimension of the Hamiltonian, and for each spin, we aim to recover the parallel and perpendicular components ($A_{\parallel},A_{\perp}$) of the hyperfine coupling tensor, which characterizes the coupling strength between the nuclei and the NV center.
@@ 4.2 Likelihood model and priors
$\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathcal{H_{k}(\mathbf{a_{k}}}),\mathbf{b},\mathbf{c}))$ $=\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}(\mathbf{d}_{i}-f_{k}(\mathcal{H}_{k}(\mathbf{A{\parallel}},\mathbf{A{\perp}},B_{z}),\lambda,N,\tau_{i}))^{2}\right),$ (13)
where $\sigma^{2}$ is a hyperparameter that relates to assumptions on the variance of the noise variance. The partition of parameters is $\mathbf{a}_{k}=(A_{\parallel,1},A_{\perp,1},A_{\parallel,2},A_{\perp,2},\dots,A_{\parallel,k},A_{\perp,k})$, $\mathbf{b}=(\lambda)$, and $\mathbf{c}=(N,\tau_{j},B_{z})$. The set $\mathbf{a}_{k}$ corresponds to the hyperfine parameters of the nuclear spins in the bath; since we do not know a priori how many nuclear spins are present, we do not know a priori how many hyperfine parameters $(A_{\parallel},A_{\perp})$ we need to recover. The continuous parameter $\mathbf{b}=(\lambda)$ that we restrict to the interval $[0,1]$ accounts for relative decoherence effects. This parameter is crucial for experimental consistency and likely reflects interactions between the NV center and unaccounted lattice impurities, such as nitrogen substitutional defects, and the broader, distant spin bath, which causes and overall decoherence decay. The known parameters $\mathbf{c}=(N,B_{z},\tau_{j})$ are the externally applied magnetic field $B_{z}$, the number of laser pulses $N$ required in the experimental protocol to obtain a coherence signal, and the amount of time $\tau_{j}$ in between sequential laser pulses.
@@ 5.2 Computational details
For each simulation, we generated coherence signals based on a set of simulated 13C that determine the number of spins and the hyperfine couplings for specified experimental parameters and relative decoherence values. We hybridize the algorithms discussed in Sec. 3 as outlined in Alg. 4, and we initialized 5 ensembles and let each ensemble evolve for 25,000 steps, and we consider the first 10,000 steps to be burn-in time. We cycle between 50 steps of RJMCMC to determine the number of spins in the system, 100 steps of PT to obtain a better fit of the hyperfine coupling Hamiltonian parameters for a fixed number of spins, and 25 steps of RWMH for the relative decoherence parameter $\lambda$.
For the lattice sites and sets of hyperfine parameters computed at each lattice site, we use data from [34] and define a cutoff of lattice sites by considering only lattice sites with hyperfine couplings with either component of $A_{\parallel}$ or $A_{\perp}$ greater than 5 kHz. For RJMCMC, we consider a maximum of $k_{\text{max}}=50$ nuclear spins to be recovered. For PT, we use $J=10$ strands and inverse temperatures $\beta_{j}=2^{-j}$. The discrete random walk over lattice sites uses a walker radius of $R=5$, and the continuous RWMH for $D$ over $[0,1]$ with reflected boundary conditions uses a walker radius of $R=0.01$. Unless stated otherwise, the default values for the simulated experiments are 250 uniformly sampled $\tau_{j}$ values between 0 and 8 $\mu$s, $N=16$ dynamical decoupling pulses, an external applied magnetic field of 311 G, and $\epsilon\sim\mathcal{N}(0,0.001)$ noise added to each data point of the simulated coherence signal. For details on how hyperparameters were chosen and how they impact the performance of the hybrid algorithm, see Appendix A.
@@
We measured the recovery rate of hyperfine parameters by examining the fraction of simulated spins that appeared in the posterior distribution across the entire spin configuration space. A recovery rate of 1 indicates that a given spin appears in every spin configuration, while a rate of 0 signifies that the spin does not appear in any posterior configuration, and we properly account for duplicate and triplicate hyperfine couplings. Additionally, we assessed the accuracy of recovering the number of spins by comparing the modes of the posterior distribution with the number of simulated spins in each hyperfine coupling grouping. We report the discrepancy in the number of spins as |mode of posterior distribution of $k$ number of simulated spins|, so a discrepancy of 0 indicates that the mode of the posterior distribution matches the number of spins that were simulated, and the greater the discrepancy the worse the hybrid MCMC procedure selects the correct Hamiltonian model dimension. We quantified false positive rate by considering the modal spin distribution corresponding to the modal model dimension of the posterior distribution as the spins that are recovered from a given coherence signal. All results were grouped by hyperfine coupling magnitude, which we define as $\sqrt{A_{\parallel}^{2}+A_{\perp}^{2}}$, and averaged 16 different simulated spin-bath configurations containing between 5 and 20 simulated spins.
@@
1:Input: data $\mathbf{d}_{d}$, total number of steps $N_{\text{tot}}$, number of RWMH steps $N_{\text{RWMH}}$, number of RJMCMC steps $N_{\text{RJMCMC}}$, number of PT steps $N_{\text{PT}}$, discrete parameter corresponding to nuclear spins $\mathbf{a}_{k}=\{p_{d_{1}},\dots,p_{d_{k}},k\}$, continuous parameters corresponding to relative decoherence $\mathbf{b}=\{\lambda\}$, fixed parameters in the forward model $\mathbf{c}=\{N,\tau,B_{z}\}$, fixed Hamiltonian model $\mathcal{H}_{k}(\mathbf{a}_{k})$ indexed by $k$, likelihood model $\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{a_{k}},\mathbf{b},\mathbf{c}))$, and hyperparameters for the individual MCMC algorithms
2: for $n=0,\dots N_{\text{tot}}-1$ do for $n=0,\dots N_{\text{tot}}-1$ do
3:      for $n=0,\dots N_{\text{tot}}-1$ do $n_{\text{RWMH}}\leftarrow 0$ $n_{\text{RWMH}}\leftarrow 0$
4:      $n_{\text{RWMH}}\leftarrow 0$ $n_{\text{RJMCMC}}\leftarrow 0$ $n_{\text{RJMCMC}}\leftarrow 0$
5:      $n_{\text{RJMCMC}}\leftarrow 0$ $n_{\text{PT}}\leftarrow 0$ $n_{\text{PT}}\leftarrow 0$
6:      $n_{\text{PT}}\leftarrow 0$ for $n_{\text{RWMH}}=0,\dots,N_{\text{RWMH}}-1$ do for $n_{\text{RWMH}}=0,\dots,N_{\text{RWMH}}-1$ do
7:          for $n_{\text{RWMH}}=0,\dots,N_{\text{RWMH}}-1$ do RWMH (Alg. 1) with 8:Input: $\mathbf{p}=\{\lambda\}$, $\mathbf{q}=\{\mathbf{a}_{k},k\},\mathbf{p}^{(0)}=\{\lambda^{n+n_{\text{RWMH}}-1}\},\mathbf{q}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RWMH}}-1},k^{n+n_{\text{RWMH}}-1}\}$ 9:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\mathbf{a}_{k}^{n},k^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$ RWMH (Alg. 1) with 8:Input: $\mathbf{p}=\{\lambda\}$, $\mathbf{q}=\{\mathbf{a}_{k},k\},\mathbf{p}^{(0)}=\{\lambda^{n+n_{\text{RWMH}}-1}\},\mathbf{q}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RWMH}}-1},k^{n+n_{\text{RWMH}}-1}\}$ 9:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\mathbf{a}_{k}^{n},k^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$
10:          RWMH (Alg. 1) with 8:Input: $\mathbf{p}=\{\lambda\}$, $\mathbf{q}=\{\mathbf{a}_{k},k\},\mathbf{p}^{(0)}=\{\lambda^{n+n_{\text{RWMH}}-1}\},\mathbf{q}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RWMH}}-1},k^{n+n_{\text{RWMH}}-1}\}$ 9:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\mathbf{a}_{k}^{n},k^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$ $n\leftarrow n+n_{\text{RWMH}}$ $n\leftarrow n+n_{\text{RWMH}}$ $n\leftarrow n+n_{\text{RWMH}}$
11:      $n\leftarrow n+n_{\text{RWMH}}$ for $n_{\text{RJMCMC}}=0,\dots,N_{\text{RJMCMC}}-1$ do for $n_{\text{RJMCMC}}=0,\dots,N_{\text{RJMCMC}}-1$ do
12:          for $n_{\text{RJMCMC}}=0,\dots,N_{\text{RJMCMC}}-1$ do RJMCMC (Alg. 2) with 13:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RJMCMC}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{RJMCMC}}-1}\}$ 14:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{k^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$ RJMCMC (Alg. 2) with 13:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RJMCMC}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{RJMCMC}}-1}\}$ 14:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{k^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$
15:          RJMCMC (Alg. 2) with 13:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RJMCMC}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{RJMCMC}}-1}\}$ 14:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{k^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$ $n\leftarrow n+n_{\text{RJMCMC}}$ $n\leftarrow n+n_{\text{RJMCMC}}$ $n\leftarrow n+n_{\text{RJMCMC}}$
16:      $n\leftarrow n+n_{\text{RJMCMC}}$ for $n_{\text{PT}}=0,\dots,N_{\text{PT}}-1$ do for $n_{\text{PT}}=0,\dots,N_{\text{PT}}-1$ do
17:          for $n_{\text{PT}}=0,\dots,N_{\text{PT}}-1$ do PT (Alg. 3) with 18:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda,k\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{PT}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{PT}}-1},k^{n+n_{\text{PT}}-1}\}$ 19:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{PT}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\lambda^{n},k^{n}\}_{n=n}^{n+N_{\text{PT}}}$ PT (Alg. 3) with 18:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda,k\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{PT}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{PT}}-1},k^{n+n_{\text{PT}}-1}\}$ 19:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{PT}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\lambda^{n},k^{n}\}_{n=n}^{n+N_{\text{PT}}}$
20:          PT (Alg. 3) with 18:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda,k\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{PT}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{PT}}-1},k^{n+n_{\text{PT}}-1}\}$ 19:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{PT}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\lambda^{n},k^{n}\}_{n=n}^{n+N_{\text{PT}}}$ $n\leftarrow n+n_{\text{PT}}$ $n\leftarrow n+n_{\text{PT}}$ $n\leftarrow n+n_{\text{PT}}$

### v2

@@ 3.4 Hybrid algorithm
Hybridizing Markov chain Monte Carlo (MCMC) algorithms can enhance sampling efficiency and model selection for complex inverse problems. A robust approach combines the random walk Metropolis-Hastings (RWMH) algorithm for exploring continuous parameter spaces with reversible jump MCMC (RJMCMC) for trans-dimensional moves that enable discrete model selection. To improve within-model fits, particularly in highly non-linear and non-convex likelihood landscapes, parallel tempering (PT) based on RWMH can be incorporated for a fixed number of discrete parameters, allowing for enhanced exploration of local minima and better posterior sampling. This multi-tiered strategy balances global model selection with efficient local exploration, mitigating issues like poor mixing and mode trapping, which are common in high-dimensional and multi-modal inverse problems.
@@ 4.1 Hamiltonian and forward models
$\mathcal{H}_{k}=D\mathcal{S}_{z}^{2}+\gamma_{e}B_{z}\mathcal{S}_{z}+\sum_{i=1}^{k}\gamma_{n}B_{z}\mathcal{I}_{z,i}+\sum_{i=1}^{k}\mathcal{S}_{z}(A_{\parallel,i}\mathcal{I}_{z,i}+A_{\perp,i}\mathcal{I}_{x,i})$ (3)
[…]
@@ 4.2 Likelihood model and priors
$\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathcal{H}_{k}(\mathbf{a_{k}}),\mathbf{b},\mathbf{c}))$ $=\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}(\mathbf{d}_{i}-f_{k}(\mathcal{H}_{k}(\mathbf{A_{\parallel}},\mathbf{A_{\perp}},B_{z}),\lambda,N,\tau_{i}))^{2}\right),$ (13)
where $\sigma^{2}$ is a hyperparameter that relates to assumptions on the variance of the noise. The partition of parameters is $\mathbf{a}_{k}=(A_{\parallel,1},A_{\perp,1},A_{\parallel,2},A_{\perp,2},\dots,A_{\parallel,k},A_{\perp,k})$, $\mathbf{b}=(\lambda)$, and $\mathbf{c}=(N,\tau_{j},B_{z})$. The set $\mathbf{a}_{k}$ corresponds to the hyperfine parameters of the nuclear spins in the bath; since we do not know a priori how many nuclear spins are present, we do not know a priori how many hyperfine parameters $(A_{\parallel},A_{\perp})$ we need to recover. The continuous parameter $\mathbf{b}=(\lambda)$ that we restrict to the interval $[0,1]$ accounts for relative decoherence effects. This parameter is crucial for experimental consistency and likely reflects interactions between the NV center and unaccounted lattice impurities, such as nitrogen substitutional defects, and the broader, distant spin bath, which causes and overall decoherence decay. The known parameters $\mathbf{c}=(N,B_{z},\tau_{j})$ are the externally applied magnetic field $B_{z}$, the number of laser pulses $N$ required in the experimental protocol to obtain a coherence signal, and the amount of time $\tau_{j}$ in between sequential laser pulses.
@@ 5.2 Computational details
For each simulation, we generate coherence signals based on a set of simulated 13C that determine the number of spins and the hyperfine couplings for specified experimental parameters and relative decoherence values. We hybridize the algorithms discussed in Sec. 3 as outlined in Alg. 4, and we initialize 5 ensembles and let each ensemble evolve for 25,000 steps, and we consider the first 10,000 steps to be burn-in time. We cycle between 50 steps of RJMCMC to determine the number of spins in the system, 100 steps of PT to obtain a better fit of the hyperfine coupling Hamiltonian parameters for a fixed number of spins, and 25 steps of RWMH for the relative decoherence parameter $\lambda$. For more details and practical strategies to choose and tune hyperparameters, see Appendix A.
[…]
We measure the recovery rate of hyperfine parameters by examining the fraction of simulated spins that appeared in the posterior distribution across the entire spin configuration space. A recovery rate of 1 indicates that a given spin appears in every spin configuration, while a rate of 0 signifies that the spin does not appear in any posterior configuration, and we properly account for duplicate and triplicate hyperfine couplings. In the simulations, where both lattice positions and hyperfine couplings are known, we impose a maximum allowed mismatch of $|\Delta A_{\parallel}|,|\Delta A_{\perp}|\leq 0.1$ kHz for a nuclear spin to be detected. This tolerance is not intended to relax the hyperfine constraint, but rather to account for small numerical errors in the DFT-computed hyperfine couplings between symmetry-related lattice sites. More precisely, let $B^{(j)}$ denote the $j$-th posterior spin-bath sample, with $j=1,\dots,M$, and define an indicator function $\mathbb{I}_{i}^{(j)}$ for the presence of a given simulated (or experimentally-confirmed) spin $s_{i}\in S$ posterior sample $B^{(j)}$, where $\mathbb{I}_{i}^{(j)}$ is 1 if $s_{i}\in B^{(j)}$ and 0 otherwise. Here, spin identity is determined by matching hyperfine parameters within the mismatch tolerance. The detection rate for an individual spin $s_{i}$ is defined as
$R_{i}=\frac{1}{M}\sum_{j=1}^{M}\mathbb{I}_{i}^{(j)},$ (14)
and the overall detection rate for a given simulated (or experimentally-confirmed) spin-bath $S$ of size $n$ is
$R=\frac{1}{n}\sum_{i=1}^{n}R_{i}.$ (15)
We then define the overall recovery rate by averaging these individual rates across all spins in the bath. When multiple simulated baths are considered (e.g., with different bath sizes), the reported values are further averaged across all simulated baths.
@@
Additionally, we assess the accuracy of recovering the number of spins by comparing the modes of the posterior distribution with the number of simulated spins in each hyperfine coupling grouping. We report the discrepancy in the number of selected spins as |(mode of posterior distribution of k)(#k)-(\# of simulated spins)|, so a discrepancy of 0 indicates that the mode of the posterior distribution matches the number of spins that are simulated, and a discrepancy greater than 0 indicates that the Hamiltonian model dimension produced by the hybrid MCMC procedure is different from the number of simulated spins. We quantify false positive rate by considering the modal spin distribution corresponding to the modal model dimension of the posterior distribution. Let $S^{*}$ denote the set of spins in the modal spin-bath configuration of size $n$, and let $B^{(j)}$ denote the $j$-th posterior spin-bath, with $j=1,\dots,M$. We define a sample-averaged false absence rate as
$FP=\frac{1}{nM}\sum_{s\in S^{*}}\sum_{j=1}^{M}\mathbb{I}_{s\notin B^{(j)}}$ (16)
where $\mathbb{I}_{s\notin B^{(j)}}$ is a non-occurrence indicator function for the presence of a spin $s$ in nuclear spin bath $B^{(j)}$. This sampled-averaged false absence rate measures the fraction of posterior samples in which spins belonging to the modal configuration are absent. We note that this is a sample-averaged measure rather than a strict false-positive rate in the classical classification sense. All results are grouped by hyperfine coupling magnitude, which we define as $\sqrt{A_{\parallel}^{2}+A_{\perp}^{2}}$, and we average across 16 different simulated spin-bath configurations containing between 5 and 20 simulated spins.
@@
1:Input: Data $\mathbf{d}_{d}$, total number of steps $N_{\text{tot}}$, number of RWMH steps $N_{\text{RWMH}}$, number of RJMCMC steps $N_{\text{RJMCMC}}$, number of PT steps $N_{\text{PT}}$, discrete parameter corresponding to nuclear spins $\mathbf{a}_{k}=\{p_{d_{1}},\dots,p_{d_{k}},k\}$, continuous parameters corresponding to relative decoherence $\mathbf{b}=\{\lambda\}$, fixed parameters in the forward model $\mathbf{c}=\{N,\tau,B_{z}\}$, family of Hamiltonians $\{\mathcal{H}_{k}\}_{k=1}^{k_{\text{max}}}$, likelihood model $\mathcal{L}(\mathbf{d}_{d}|f_{k}(\mathbf{a_{k}},\mathbf{b},\mathbf{c}))$, and hyperparameters for the individual MCMC algorithms
2: for $n=0,\dots N_{\text{tot}}-1$ do
3:   $n_{\text{RWMH}}\leftarrow 0$
4:   $n_{\text{RJMCMC}}\leftarrow 0$
5:   $n_{\text{PT}}\leftarrow 0$
6:   for $n_{\text{RWMH}}=0,\dots,N_{\text{RWMH}}-1$ do
7:    RWMH (Alg. 1) with 8:Input: $\mathbf{p}=\{\lambda\}$, $\mathbf{q}=\{\mathbf{a}_{k},k\},\mathbf{p}^{(0)}=\{\lambda^{n+n_{\text{RWMH}}-1}\},\mathbf{q}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RWMH}}-1},k^{n+n_{\text{RWMH}}-1}\}$ 9:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RWMH}}}=\{\mathbf{a}_{k}^{n},k^{n}\}_{n=n}^{n+N_{\text{RWMH}}}$
10:    $n\leftarrow n+n_{\text{RWMH}}$
11:   for $n_{\text{RJMCMC}}=0,\dots,N_{\text{RJMCMC}}-1$ do
12:    RJMCMC (Alg. 2) with 13:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{RJMCMC}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{RJMCMC}}-1}\}$ 14:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}=\{\lambda^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$, $\{k^{n}\}_{n=n}^{n+N_{\text{RJMCMC}}}$
15:    $n\leftarrow n+n_{\text{RJMCMC}}$
16:   for $n_{\text{PT}}=0,\dots,N_{\text{PT}}-1$ do
17:    PT (Alg. 3) with 18:Input: $\mathbf{p}=\{\mathbf{a}_{k}\}$, $\mathbf{q}=\{\lambda,k\},\mathbf{p}^{(0)}=\{\mathbf{a}_{k}^{n+n_{\text{PT}}-1}\},\mathbf{q}^{(0)}=\{\lambda^{n+n_{\text{PT}}-1},k^{n+n_{\text{PT}}-1}\}$ 19:Output: $\{\mathbf{p}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\mathbf{a}_{k}^{n}\}_{n=n}^{n+N_{\text{PT}}}$, $\{\mathbf{q}^{n}\}_{n=n}^{n+N_{\text{PT}}}=\{\lambda^{n},k^{n}\}_{n=n}^{n+N_{\text{PT}}}$
20:    $n\leftarrow n+n_{\text{PT}}$